# Kareem Amin  Alex Bie  Weiwei Kong  Alexey Kurakin Natalia Ponomareva  Umar Syed  Andreas Terzis  Sergei Vassilvitskii Google

### v1

@@ B.1 Logits clipping function
Appendix C Proof of Theorem 1
Our proof of Theorem 1 is organized into sections. §C.1 provides basic definitions. §C.2 and §C.3 establish key results related to composition and sensitivity. §C.4 proves the privacy of simpler mechanisms that each account for a portion of the functionality of Algorithm 1. C.5 puts all the pieces together and completes the proof.
C.1 Definitions
@@ C.2 Composition
Parallel composition is a well-known technique in differential privacy that is useful for establishing privacy guarantees in scenarios where a mechanism is independently applied to disjoint subsets of a dataset. Many versions of parallel composition require that the subsets are chosen in a fully data-independent manner. We show that the same result holds under a weaker assumption.
@@
Let $k$ be a positive integer. Let $f$ be a function that maps prompts into $[k]$. For any dataset of prompts $D$ and $i\in[k]$ let
@@
Let $M$ be a mechanism that satisfies $\rho$-zCDP. If $\widehat{M}$ is the mechanism defined by
@@
$D_{\alpha}(\widehat{M}(D)\parallel\widehat{M}(D^{\prime}))$ $=\sum_{i=1}^{k}D_{\alpha}(M(D_{i})\parallel M(D^{\prime}_{i}))$
$=D_{\alpha}(M(D_{j})\parallel M(D^{\prime}_{j}))$
@@ C.3 Sensitivity analysis
$d(Z,\mathbf{z})=\left\lVert\frac{1}{s}\sum_{\mathbf{z}^{\prime}\in Z}p_{\mathbf{z}^{\prime}}-p_{\mathbf{z}}\right\rVert_{1}$
@@
The function $\ell$ has sensitivity $\frac{c}{s}$, and for all $\mathbf{z}\in\mathbb{R}^{v}$, the function $d(\cdot,\mathbf{z})$ has sensitivity $\frac{1}{s}$.
@@
$\left|d(Z,\mathbf{z})-d(D^{\prime},\mathbf{z})\right|$
@@ C.4 Constituent mechanisms
Consider a function $f:\mathcal{D}\rightarrow\mathbb{R}^{v}$ with sensitivity $\Delta$. By an analysis of the exponential mechanism due to Cesar and Rogers [2021], [^6] See also Rogers and Steinke [2021]. choosing a token according to the distribution $\operatorname{softmax}(\frac{\varepsilon}{2\Delta})$ satisfies $\frac{1}{8}\varepsilon^{2}$-zCDP. Observe that mechanism $M$ is the exponential mechanism with $f=\frac{1}{\tau}\ell$, which by Lemma 3 has sensitivity $\frac{c}{s\tau}$. ∎
@@
8:     $\hat{d}\leftarrow d(Z,\mathbf{z}_{\operatorname{public}})+\textrm{Laplace}(2\sigma)$
9:     if $\hat{q}>\hat{\theta}$ then
@@
Observe that mechanism $M$ is an instance of the AboveThrehold mechanism [Dwork et al., 2009], which accepts a private dataset, a threshold, and a sequence of queries as input. In each iteration, the AboveThreshold mechanism applies the next query in the sequence to the dataset and compares it to a noisy threshold, and returns the index of the first query that exceeds the threshold. The queries can be chosen adaptively and adversarially. In mechanism $M$, each query is specified by a token sequence $\mathbf{x}$, and the index of the first query that exceeds the threshold is determined by the length of the returned token sequence. Furthermore, by Lemma 3 each query has sensitivity $\frac{1}{s}$. Thus by the analysis due to Dwork et al. [2009], mechanism $M$ satisfies $\frac{2}{s\sigma}$-differential privacy, which by Bun and Steinke [2016] implies that mechanism $M$ satisfies $\frac{2}{(s\sigma)^{2}}$-zCDP. ∎
C.5 Putting it all together
Consider a sequence of iterations of the inner loop of Algorithm 1 in which the value of $t$ (the private token counter) is constant. Observe that the operation of Algorithm 1 during these iterations is equivalent to the sequential composition of Algorithms 2 and 3, since these iterations generate zero or more public tokens followed by a private token. [^7] The special treatment of the <eos> token complicates this story a little, but we can always assume that the LLM ignores any tokens before the last <eos> token. Moreover, there are at most $r$ such sequences of iterations, since $r$ is an upper bound on the private token counter for any batch. By Lemmas 1, 4 and 5 we have that Algorithm 1 applied to a single batch satisfies $\rho$-zCDP (where $\rho$ is specified in the statement of Theorem 1). And therefore by Assumption 1 and Lemma 2 we have that Algorithm 1 applied to the whole dataset satisfies $\rho$-zCDP. It remains to convert this zCDP guarantee to an $(\varepsilon,\delta)$-differential privacy guarantee, which we do two different ways using two existing results: Corollary 13 due to Canonne et al. [2020] and Lemma 3.5 due to Bun and Steinke [2016].
Appendix D Privacy-equivalent Gaussian noise
Given the average token distribution $\bar{\mathbf{p}}$ in a batch, Tang et al. [2024] protect the privacy of $\bar{\mathbf{p}}$ by using the Gaussian mechanism, which achieves $(\varepsilon,\delta)$-differential privacy with $\varepsilon=\frac{\sqrt{2\log(1.25/\delta)}}{s\sigma}$, where $s$ is the batch size and $\sigma$ is the standard deviation of the noise added to each probability in $\bar{\mathbf{p}}$. On the other hand, we use the exponential mechanism to protect the privacy of a sample drawn from $\bar{\mathbf{p}}$, which achieves $\varepsilon$-differential privacy with $\varepsilon=\frac{2c}{s\tau}$, where $c$ is the maximum absolute value of any log-probability in the batch and $\tau$ is the sampling temperature.
@@ Appendix D Privacy-equivalent Gaussian noise
The analysis in Theorem 8 of Balle and Wang [2018] does not admit a closed-form solution. Instead, we binary search for a solution to:
@@
Finally, equating the zCDP loss for the exponential mechanism given by $\frac{\varepsilon^{2}}{8}=\frac{c^{2}}{2s^{2}\tau^{2}}$ (Cesar and Rogers [2021]) to that of the Gaussian mechanism given by $\frac{1}{2s^{2}\sigma^{2}}$ (Bun and Steinke [2016]), yields $\sigma=0.2$.
Appendix E Experiment details
E.1 Hyperparameter tuning
@@
Based on initial experiments, we found that setting $c=10$ and $\tau=2$ produced well formed text, so we fix $c=10$ and try a low temperature ($\tau=1.5$) and a high temperature ($\tau=2.25$) setting. At $\tau=2.25$, we observed text degeneration. This is due to the combination of Gemma’s large vocabulary (256K) and clipping, which raises the “probability floor” of nonsense tokens. So for $\tau=2.25$ settings only, we follow Tang et al. [2024] and reduce the vocabulary to the public prediction’s top 1024. We emphasize that (1) we do not do this for any of the other settings of $\tau$, and (2) use a larger value than prior work (they use top 100).
@@ E.2 Prompts used
Appendix F Artifacts
Tables 1(a) and 1(b) list all artifacts we use in this work. AGNews, TREC, DBPedia, MIT-G, MIT-D, IMDB, and Yelp are all standard academic datasets permissible for research use; we cite their original publications when introduced. WikiMoviesJSON is scraped from Wikipedia data, courtesy of [Rust, 2024]; their work is covered by an MIT license. Wikipedia content is licensed under the Creative Commons Attribution-ShareAlike 4.0 International License (CC BY-SA) and the GNU Free Documentation License (GFDL).
We use open-source models BERT-Base, released by [Turc et al., 2019], and Gemma. Our use of Gemma for academic purposes is in accordance of the Gemma terms of use: https://ai.google.dev/gemma/terms. GPT-3 is accessible for academic purposes under OpenAI’s terms of use, which supports educational and research activities. LaMDA 8B is not publically available, but we received sufficient authorization to use it for the academic purposes of this paper.
Appendix G Compute budget
All experiments for synthetic data generation run on Gemma 2B 1.1 IT. A run of synthetic data generation takes between 8-48 accelerator hours. Including exploratory runs and hyperparameter search, the total compute budget for this project is roughly 14,000 accelerator hours.

### v2

@@ Discussion.
Appendix D Proof of Theorem 1
Our proof of Theorem 1 is organized into sections. §D.1 provides basic definitions. §D.2 and §D.3 establish key results related to composition and sensitivity. §D.4 proves the privacy of simpler mechanisms that each account for a portion of the functionality of Algorithm 1. §D.5 puts all the pieces together and completes the proof.
D.1 Definitions
@@ D.2 Composition
Parallel composition is a well-known technique in differential privacy that is useful for establishing privacy guarantees in scenarios where a mechanism is independently applied to disjoint batches of a dataset. Many versions of parallel composition require that the batches are chosen in a fully data-independent manner. We show that the same result holds under a weaker assumption.
@@
Let $k>1$ be the number of batches. Let $f$ be a batch assignment function that satisfies Assumption 1. In other words, let $f$ be a function that map each prompt to and element of $[k]$. For any dataset of prompts $D$ and $i\in[k]$ let
@@
be the $i$th batch of prompts. Let $M$ be a mechanism that satisfies $\rho$-zCDP. If $\widehat{M}$ is the mechanism defined by
@@
$D_{\alpha}(\widehat{M}(D)\|\widehat{M}(D^{\prime}))$ $=\sum_{i=1}^{k}D_{\alpha}(M(D_{i})\|M(D^{\prime}_{i}))$
$=D_{\alpha}(M(D_{j})\|M(D^{\prime}_{j}))$
@@ D.3 Sensitivity analysis
$\mathrm{d}(Z,\mathbf{z})=\left\lVert\frac{1}{s}\sum_{\mathbf{z}^{\prime}\in Z}p_{\mathbf{z}^{\prime}}-p_{\mathbf{z}}\right\rVert_{1}$
@@
The function $\ell$ has sensitivity $\frac{c}{s}$, and for all $\mathbf{z}\in\mathbb{R}^{v}$, the function $\mathrm{d}(\cdot,\mathbf{z})$ has sensitivity $\frac{1}{s}$.
@@
$\left|\mathrm{d}(Z,\mathbf{z})-d(D^{\prime},\mathbf{z})\right|$
@@ D.4 Constituent mechanisms
Consider a function $f:\mathcal{D}\rightarrow\mathbb{R}^{v}$ with sensitivity $\Delta$. By an analysis of the exponential mechanism due to Cesar and Rogers (2021), [^6] See also Rogers and Steinke (2021). choosing a token according to the distribution $\operatorname{softmax}(\frac{\varepsilon}{2\Delta})$ satisfies $\frac{1}{8}\varepsilon^{2}$-zCDP. Observe that mechanism $M$ is the exponential mechanism with $f=\frac{1}{\tau}\ell$, which by Lemma 3 has sensitivity $\frac{c}{s\tau}$. ∎
@@
8:     $\hat{d}\leftarrow\mathrm{d}(Z,\mathbf{z}_{\operatorname{public}})+\textrm{Laplace}(2\sigma)$
9:     if $\hat{d}\geq\hat{\theta}$ then
@@
Observe that mechanism $M$ is an instance of the AboveThrehold mechanism (Dwork et al., 2009), which accepts a private dataset, a threshold, and a sequence of queries as input. In each iteration, the AboveThreshold mechanism applies the next query in the sequence to the dataset and compares it to a noisy threshold, and returns the index of the first query that exceeds the threshold. The queries can be chosen adaptively and adversarially. In mechanism $M$, each query is specified by a token sequence $\mathbf{x}$, and the index of the first query that exceeds the threshold is determined by the length of the returned token sequence. Furthermore, by Lemma 3 each query has sensitivity $\frac{1}{s}$. Thus by the analysis due to Dwork et al. (2009), mechanism $M$ satisfies $\frac{2}{s\sigma}$-differential privacy, which by Bun and Steinke (2016) implies that mechanism $M$ satisfies $\frac{2}{(s\sigma)^{2}}$-zCDP. ∎
D.5 Putting it all together
Consider a sequence of iterations of the inner loop of Algorithm 1 in which the value of $t$ (the private token counter) is constant. Observe that the operation of Algorithm 1 during these iterations is equivalent to the sequential composition of Algorithms 2 and 3, since these iterations generate zero or more public tokens followed by a private token. [^7] The special treatment of the <eos> token complicates this story a little, but we can always assume that the LLM ignores any tokens before the last <eos> token. Moreover, there are at most $r$ such sequences of iterations, since $r$ is an upper bound on the private token counter for any batch. By Lemmas 1, 4 and 5 we have that Algorithm 1 applied to a single batch satisfies $\rho$-zCDP (where $\rho$ is specified in the statement of Theorem 1). And therefore by Assumption 1 and Lemma 2 we have that Algorithm 1 applied to the whole dataset satisfies $\rho$-zCDP. It remains to convert this zCDP guarantee to an $(\varepsilon,\delta)$-differential privacy guarantee, which we do two different ways using two existing results: Corollary 13 due to Canonne et al. (2020) and Lemma 3.5 due to Bun and Steinke (2016).
Appendix E Privacy-equivalent Gaussian noise
Given the average token distribution $\bar{\mathbf{p}}$ in a batch, Tang et al. (2024) protect the privacy of $\bar{\mathbf{p}}$ by using the Gaussian mechanism, which achieves $(\varepsilon,\delta)$-differential privacy with $\varepsilon=\frac{\sqrt{2\log(1.25/\delta)}}{s\sigma}$, where $s$ is the batch size and $\sigma$ is the standard deviation of the noise added to each probability in $\bar{\mathbf{p}}$. On the other hand, we use the exponential mechanism to protect the privacy of a sample drawn from $\bar{\mathbf{p}}$, which achieves $\varepsilon$-differential privacy with $\varepsilon=\frac{2c}{s\tau}$, where $c$ is the maximum absolute value of any log-probability in the batch and $\tau$ is the sampling temperature.
@@ Appendix E Privacy-equivalent Gaussian noise
The analysis in Theorem 8 of Balle and Wang (2018) does not admit a closed-form solution. Instead, we binary search for a solution to:
@@
Finally, equating the zCDP loss for the exponential mechanism given by $\frac{\varepsilon^{2}}{8}=\frac{c^{2}}{2s^{2}\tau^{2}}$ (Cesar and Rogers (2021)) to that of the Gaussian mechanism given by $\frac{1}{2s^{2}\sigma^{2}}$ (Bun and Steinke (2016)), yields $\sigma=0.2$.
Appendix F Experiment details
F.1 Hyperparameter tuning
@@
Based on initial experiments, we found that setting $c=10$ and $\tau=2$ produced well formed text, so we fix $c=10$ and try a low temperature ($\tau=1.5$) and a high temperature ($\tau=2.25$) setting. At $\tau=2.25$, we observed text degeneration. This is due to the combination of Gemma’s large vocabulary (256K) and clipping, which raises the “probability floor” of nonsense tokens. So for $\tau=2.25$ settings only, we follow Tang et al. (2024) and reduce the vocabulary to the public prediction’s top 1024. We emphasize that (1) we do not do this for any of the other settings of $\tau$, and (2) use a larger value than prior work (they use top 100).
@@ F.2 Prompts used
Appendix G Artifacts
Tables 1(a) and 1(b) list all artifacts we use in this work. AGNews, TREC, DBPedia, MIT-G, MIT-D, IMDB, and Yelp are all standard academic datasets permissible for research use; we cite their original publications when introduced. WikiMoviesJSON is scraped from Wikipedia data, courtesy of (Rust, 2024); their work is covered by an MIT license. Wikipedia content is licensed under the Creative Commons Attribution-ShareAlike 4.0 International License (CC BY-SA) and the GNU Free Documentation License (GFDL).
We use open-source models BERT-Base, released by (Turc et al., 2019), and Gemma. Our use of Gemma for academic purposes is in accordance of the Gemma terms of use: https://ai.google.dev/gemma/terms. GPT-3 is accessible for academic purposes under OpenAI’s terms of use, which supports educational and research activities. LaMDA 8B is not publically available, but we received sufficient authorization to use it for the academic purposes of this paper.
Appendix H Compute budget
Our main experiments for synthetic data generation run on Gemma 1.1 2B IT. A run of synthetic data generation takes between 8-48 accelerator hours. Including exploratory runs and hyperparameter search, the total compute budget for this project is roughly 14,000 accelerator hours.